jack usually mows the lawn in 3 hours. Sally mows the same yard in 4 hours. How much time would it take for them to mow the lawn together?
Found 2 solutions by richwmiller, solver91311:
Answer by richwmiller(17219)

x/3+x/4=1
4x+3x=12
7x=12
x=12/7=1.71429 hours

Answer by solver91311(24713)



If A can do a job in x time periods, then A can do of the job in 1 time period. Likewise, if B can do the same job in y time periods, then B can do of the job in 1 time period.

So, working together, they can do



of the job in 1 time period.

Therefore, they can do the whole job in:



time periods.


John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism